OpenStudy (anonymous):

4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

you have to distribute

OpenStudy (anonymous):

4(x - 6) = -2(x - 3) Start by distributing 4 over to x and -6 4x-24=-2(x - 3) now distribute the -2 over to x and -3 4x-24=-2x+6 now move the x value's to the left side and number to right side

OpenStudy (anonymous):

The answer is 6

OpenStudy (anonymous):

You are pretty close, but the answer is not 6

OpenStudy (anonymous):

-6-18